OpenStudy (anonymous):

3u+2v= -17 2u+v = -8 The solution is There are infinitely many solutions There is no solution

OpenStudy (anonymous):

2(3u+2v=-17) 3(2u+v=-8) so, 6u+4v=-34 and 6u+3v=-24 6u=34-4v & 6u=-24-3v so, 34-4v=-24-3v 34+24=v=58 & 6u=34-198 thus u=-33. Hence there is only one solution.

OpenStudy (anonymous):

These are simultaneous equations hence only 1 solution exists....